### v1

# Diffusion Models for Computational Design at the Example of Floor Plans

Abstract
AI Image generators based on diffusion models are widely discussed recently for their capability to create images from simple text prompts. But, for practical use in civil engineering they need to be able to create specific construction plans for given constraints. Within this paper we explore the capabilities of those diffusion-based AI generators for computational design at the example of floor plans and identify their current limitation. We explain how the diffusion-models work and propose new diffusion models with improved semantic encoding. In several experiments we show that we can improve validity of generated floor plans from 6 % to 90 % and query performance for different examples. We identify short comings and derive future research challenges of those models and discuss the need to combine diffusion models with building information modelling. With this we provide key insights into the current state and future directions for diffusion models in civil engineering.
keywords:
diffusion models, deep learning, computational design, building information modelling
1 Introduction
1.1 Motivation
[…]
The new AI art generators show a remarkable capability in generating images of architectural design sketches and frontal or perspective views that display a higher degree of creativity, even if they are technically just recombining the training material. This opens the question if they can be used to more challenging tasks like generating floor plans or sectional drawings which are nowadays often manually drawn on paper or in CAD (Computer Aided Design) tools.
The process of planning constructions in civil engineering is still a very manual and time-consuming process. It possesses enormous potential for boosting productivity by 50 % to 60 %, which could result in an extra 1.6 trillion USD being added annually to the value of the construction industry barbosa2017reinventing . AI approaches are considered a core enable of this productivity boost PAN2021103517 ; pan2023integrating ; debrah2022artificial .
The recent breakthroughs of the AI art generation tools are based on the introduction of diffusion models in deep learning. They were first introduced by Sohl-Dickstein et al. in sohl2015deep and successively destroy information in an image during the training phase to learn how to recreate it. We will explain this process in more detail in Section 1.2.4.
The main topic of our paper is how these diffusion models can be used for computational design of architectural drawings like floors plans and how we can improve them to do this better. Within this paper we want to:
  • Evaluate the capability of diffusion models in computational design
  • Discuss the requirements of diffusion models for architectural planning
  • Propose different semantic image-based diffusion model encodings
  • Compare the performance and workflows of those encodings
  • Discuss if and how we can directly use BIM data for diffusion models
  • Derive the requirements and architecture for a BIM-based diffusion model
  • Identify future directions for diffusion models in computational design
1.2 State of the Art
1.2.1 Parametric Design
The idea of automating design steps with technology is an old one and various approaches developed over the time from parametric design to generative, and algorithmic design. Caetano, Santos, and Leitão combine them in their review under the term Computational Design caetano2020computational .
The idea of Parametric Design is that components are created and shaped according to parameterized functions in a computer, in contrast to being designed directly gursel2012creative ; jabi2013parametric ; schnabel2007parametric . It is founded in modelling shape attributes trough mathematical functions and was already used before computers were introduced e. g. by Antoni Gaudi in designing elements in the Sagrada Familia burry1993expiatory . Nowadays, the shape attributes are altered by computers which usually generates a large set of candidates that are then for example validated structurally monedero2000parametric ; rempling2019automatic or in their energy performancegranadeiro2013building to identify good ones to be presented to the user.
1.2.2 Generative Design
Generative Design is using generative computer approaches like evolutionary or agent-based algorithms to generate and evaluate designs caldas2008generation ; krish2011practical . More recent approaches are starting to adopt Graph representations gan2022bim and Deep Neural Networks for generative tasks min2023floor ; rahbar2022architectural ; weber2022automated . The challenge in those approaches is that they are bound in their generation capability by the underlying generation rules and representation (e. g. genome encoding) and cannot in a creative way recombine them. Further they are prone to generate random results outside of the requested targets due to their probabilistic or non-deterministic nature of their search caetano2020computational ; chaszar2016generating .
1.2.3 Algorithmic Design
Algorithmic design approaches usually describe computer programs that are used to generate using plans based on rules and logic implemented in code or code grammars humppi2016algorithm ; oxman2017thinking ; caetano2019integration . More recent approaches combine such a shape grammar with reinforcement learning to generate energy efficient designs mandow2020architectural . These approaches require understanding of the code to control the generation process, which makes them less applicable in practise.
1.2.4 Diffusion models
Diffusion Models revolutionized the area of generating art with AI. They were first introduced in sohl2015deep and are inspired by non-equilibrium thermodynamics. Diffusion-models are based on a Deep Neural Network (DNN) architecture that does two things: First, it successively destroys the information in the image it is trained on by adding white noise in a so called forward diffusion step. In each step it trains a DNN to be able to recover the lost information in a reverse diffusion step. The resulting image now contains some of the original information like style and colour and some newly added information which is associated to tags describing the image. As the DNN is only trained to repair a small piece of lost information in the image, the generation approach becomes more stable and solve a main challenge with traditional approaches that try to generate the whole image from scratch and used to fail in the process.
This destroy and repair process is then repeated at different levels of resolution. This removes the limitation that the DNN can only repair lost information as we can now create new images by first creating them at a very low resolution and then successively repairing them until we have a high-resolution image. It also allows the DNN for the different resolutions to specialize on different visual aspects. For example, lower resolutions layers learn things like image composition, middle layers learn object shapes, and high resolutions learn object surface materials and colours.
(a) Step 1
(b) Step 2
(c) Step 3
Figure 1: Examples of iterative denoising steps for a building floor plan in Midjourney v4.
This basic idea of diffusion was then used to create neural network architectures that are capable of using reverse diffusion (denoising) to create high-quality image outputs from noisy inputs ho2020denoising . The concept is nowadays widely applied beyond image generation in natural language processing, video and 3D model generation, pose detection, timeseries modelling, etc. yang2023diffusion . An example of this step-by-step process is shown in Fig. 1. Configuring these models to create the desired outputs used to be a process that required expert knowledge. More recent architecture variants like OpenAI’s GLIDE model nichol2021glide contain an encoder, which can take an arbitrary text prompt by a user and create a valid text encoding that can be fed into the connected diffusion model. This architecture also includes a second model, which upsamples the result of the diffusion model.
Most current commercial models usually use the CLIP (Contrastive Language Image Pre-training) architecture presented in radford2021learning and ramesh2021zero . CLIP is responsible for training the encoder and determining how the text encodings are linked to image parts in the diffusion model. From this point, every platform contains slight differences in model and encoder architecture. The specific encoder used there is called unClip, which includes an image encoder and encodes both text and image inputs into a joint representation space from which the diffusion model can create an image Ramesh22 .
[…]
1.2.5 Fine-tuning Diffusion Models
[…]
Fortunately, the new generation of diffusion models allows to refine these federated models with a few examples to a new domain, which is called fine-tuning or few-shot-learning. There are different approaches for doing that. One approach is to retrain the whole diffusion-model on the new dataset and refine the weights in the deep neural network (DNN) Ruiz_2023_CVPR . This creates a variant of the initial federated diffusion-model that is of the same size as the original network and quite computational expensive to refine all the weights on the DNN.
To avoid this, other approaches introduce new layers into the federated diffusion model and modify only their weights. This results in smaller and faster to train models as the original federated model is not modified and fewer weights need to be fine-tuned hu2021lora ; roich2022pivotal . The additional layers also enable the models to also learn new capabilities, like transforming outline sketches into full images zhang2023adding .
The final approach to fine-tune is not to modify the DNN, but, the input vector from the prompt. The text prompt send to the diffusion model needs to be encoded as number vector to be used as input layer for the DNN. The finetuning changes the weights of this word embedding vector. A process called textual inversion gal2022textual . The benefit of this approach is that it is very fast to train and results in very small fine-tuned model that only contains the modified word embedding vector.
1.2.6 3D Diffusion models
[…]
2 Diffusion models for Floor Plans
2.1 Integrated workflow
One goal of the paper is to understand how generative AI tools can support practitioners in their typical design workflows. For this we want to compare the workflows for the different diffusion models. For this comparison we want first want to discuss a idealized workflows would like that fully integrates and unlocks the potential of generative AI tools in the design process in our opinion. This workflow is illustrated in Fig. 2.
Figure 2: Idealized integrated workflow for AI-based design
In this workflow, a planner uses a text interface or speech to text interface to specify the keywords describing his targeted design. As alternative input the user may be able to sketch out a rough plan to provide e. g. the general shape of a more complex planning tasks. The AI system would refine these initial inputs by generating questions to collect additional requirements and clarify uncertainties in a dialog with the user. These requirements may encompass functional (e. g. room types, size, etc.), non-functional (e. g. materials), stylistic (e. g. shape and colours), or computational constraints (e. g. structural dynamics, energy-efficiency). It would then generate multiple design options using a federated diffusion model, validate and rank them and allow the user to select the best ones and edit areas that require more work. The final design should be exported to a BIM model like IFC that the user can easily import in his tools for further use. This workflow is based on the common workflows for diffusion models ec3 . We consider it also good for civil engineering, because the user has multiple ways to control the design process from specify and refining requirements to editing model parts.
Within the next sections we will first discuss qualitatively, what elements of this workflow are already feasible with the existing AI tools and then evaluate their performance in different experiments in section 3. From this we derive further research topics needed to realize this idealized workflow in section 4.
2.2 Out-of-the-box Diffusion Model $\mathcal{B}$
Many users are adopting the new AI art tools to generate architectural drawings. The study in ec3 has shown that already 7 % of all prompts on the Midjourney platform are related to architecture with 14.741 prompts in mean per day. Of those prompts 43 % are related to floor plans.
Midjourney can be used with the Discord messaging app as it runs on the servers of the provider. This on the one hand makes it easy to use, as no local model needs to be installed and configured. On the other hand, all queries are run on the companies servers and by default are visible to other users. This poses a privacy and data security risk.
Stable Diffusion is a alternative generative AI solution that can be installed and operated locally. This overhead brings the benefit to the users to execute their prompts in private and to retrain the model for their specific applications.
The out-of-the-box approach for a user testing the technology is to use the out-of-the box model to generate his floor plans. We call this approach our baseline approach $\mathcal{B}$. Fig. 3 show the workflow for this. The user writes a text query describing the design they would like to have and sends it to the generative AI. Out-of-the-box, the generative AI uses a federated Diffusion model and generates one or multiple binary images as response to the query.
The workflow of this out-of-the-box approach has multiple differences in comparison to Fig. 2. We do not have the refinement loops like that the AI model generates refining questions or the section editing parts of the generated design as they are not fully supported by the current technologies ploennigs2022ai . Further, the out-of-the-box approach creates only a binary image of a floor plan instead of directly generating a BIM model. In theory, this binary image could be used to automatically reconstruct a BIM model. However, this is a non-trivial problem still in research gimenez2016automatic ; zeng2019deep ; pizarro2022automatic . In practise, the user needs to manually create the BIM model from on the generated image.
Figure 3: Generation Workflow for out-of-the-box Diffusion Model
2.3 Fine-tuned Diffusion Model $\mathcal{R}$
The out-of-the-box approach $\mathcal{B}$ is prone to creating faulty results, as we will confirm in the experimental section 3. Fig. 4 shows an example floor plan generated by Stable Diffusion for the query “building floor plan of a one family house with a garden”. The image shows some resemblance to a floor plan, but also many incorrect elements. The reason for the low-quality result is that the out-of-the-box model is trained on a wide variety of images and will respond with images with a large variance and not the expected narrow focus on floor plans.
Figure 4: Example floor plan of a house with garden created by the out-of-the-box model $\mathcal{B}$.
Therefore, it is necessary to fine-tune the diffusion model to correctly understand what kind of floor plans we expect. This can be done by retraining the model with new sample images and queries that resemble the expected results as discussed in section 1.2.5. This modifies our workflow as illustrated in Fig. 5. We now have a training and a generation phase. In the training phase we fine-tune an existing federated diffusion model with a small set of sample images and queries. We use the tuned diffusion model in the generation phase to generate then designs that are closer to the training data.
(a) Training
(b) Generation
Figure 5: Workflow for a refined 2D/3D bitmap-based Diffusion Model
2.4 Fine-tuned Diffusion Models with Semantic Encoded rooms $\mathcal{SR}$ and elements $\mathcal{SE}$
(a) $\mathcal{R}$
(b) $\mathcal{SR}$
(c) $\mathcal{SE}$
(d) $\mathcal{SRE}$
Figure 6: Example generation results of the four retrained styles.
In previous work ec3 it was shown that one of the main reasons why diffusion models struggle with floor plans is the missing semantic understanding of the elements in the image. While a diffusion model may learn that the term ”floor plan” is correlated to rectangular lines, it does not really understand that these lines are walls enclosing rooms. The moment we add more symbols to the image like doors, windows, furniture or assets, the model will struggle understanding the meaning of these symbols and mix them up as visible in Fig. 4. The question is: Can make it easier for the diffusion model to learn these semantics?
We propose a novel approach to encode these semantics in floor plans to improve the fine-tuning of diffusion models utilizing the following three strategies:
Simplify:
The more different semantics a floor plan contains the harder it is for the diffusion model to differentiate. Thus, it is important to remove unnecessary elements (furniture, assets, etc.) and focus on the relevant semantics for floor planning (walls, doors, and windows). The removed unnecessary elements may be added in a following in-painting step in the section editing step from Fig. 2. This will be our baseline for the retraining of model $\mathcal{B}$, which will be called $\mathcal{R}$ (Reduced).
Encode:
When we look at a floor plan, we know that thick lines are walls and the enclosed white space is a room. We also understand its semantic function (kitchen, bed room, living room) from the symbols of appliances that it contains (oven/fridge, bed, couch). When we remove those symbols—to simplify the model—the diffusion model (and we) cannot understand this semantic. So, we need a way to re-encode this information, for example by filling the rooms with different colours like blue for kitchen, orange for a bed room, or green for a living room. We call this semantic colouring of rooms $\mathcal{SR}$.
Contrast:
Floor plans are often black and white (BW) and diffusion models are designed to encode colour images. Thus, we lose lots of information capacity by just training them on black and white images. We already used colours to encode room semantics in the last step. The question is if we can further simplify the floor plans by replacing the symbols for doors and windows with colours instead. We call this semantic colouring of elements $\mathcal{SE}$. The combined semantic encoding of room and elements is the $\mathcal{SRE}$ model.
2.5 Fine-tuning Methodology
In order to fine-tune Stable Diffusion to the different semantic encodings, we implemented an algorithmic design approach in Python that generates a desired number of floor plans and then encodes it in each of the four different proposed styles $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, and $\mathcal{SRE}$ as well as a description prompt for each image. This ensures that all four models are on even ground during the training phase. Fig. 6 shows example images for each of the proposed encodings.
The generated prompts describe the generated floor plan and follows the common guidelines for Stable Diffusion prompts, specifically for automatic1111: Individual concepts are separated by commas and round brackets. Commas act as conceptual separators and brackets assign prompt weights to each of the concepts (always kept at 1 here). There is also a one token long style descriptor that needs to be included at the beginning to invoke the trained style. Each concept descriptor contains a number, a quantity descriptor (few or many), the word associated with the element itself, and a colour.
Figure 7: Example prompt used in training and generation: Green - building shape; Magenta - kitchen; Cyan - bathroom; Yellow - living rooms, Blue - windows; Red - doors
The training data was created to cover various possible floor plan designs from different building shapes up to buildings that have special arrangements like having no windows or a specific number of doors. The individual symbologies (like doors or living rooms) are also trained through images that only feature these symbols in several orientations.
The following training images were generated for each style:
  • Two images each of individual doors, windows, three images each of living rooms, bathrooms and kitchens
  • Ten floor plans each in which an element has an altered colour specification (green)
  • Ten floor plans each for different building shapes: l-shaped, o-shaped, c-shaped, square, rectangle, multiple buildings
  • Ten floor plans each for negations, i. e. elements whose count is 0 and quantity descriptors is “no”
  • Ten floor plans each for the few and many cases of each kind of element
  • Ten floor plans each in which the count of an element is set to 2, 4 or 6
We fine-tuned a Stable Diffusion model v2.1 for each encoding $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, and $\mathcal{SRE}$ using Textual Inversions gal2022textual as preliminary trials showed that it offers fastest training time at equal result performance for our use case in comparison to hu2021lora and Ruiz_2023_CVPR .
3 Experiments
3.1 Validation Methodology
To compare the performance of the different un-tuned and fined-tuned diffusion models we defined a large set of 43 test prompts that evaluate different features representing practical request as well as challenging situations for the diffusion models. The prompts classify into the following categories:
Valid Plan:
First, we evaluate how many of the generated images show a valid floor plan that is not cut off and contains at least some recognizable symbology.
Overfit Check:
Second, we check if the fine-tuned models do not overfit on the training samples, by prompting for human faces expecting that we do not get floor plans.
Quantify Objects/Rooms:
To see if the model can understand quantities we query for “few” (less than 7) and “many” (more than 6) windows/doors and rooms/kittchen/bathrooms.
Count of Objects/Rooms:
Instead of asking for fuzzy quantities, we ask for specific counts (2, 4, 6) of windows/doors and rooms/kittchen/bathrooms.
Remove Objects/Rooms:
To check if the model has semantic understanding of core terms, we ask it to remove them and e. g. create a building with no windows/doors or no kitchens, etc. (Note: this is not done through negative prompts, as we want to test the ability of the models to turn structured queries into semantically sound plans.)
Recolour Objects/Rooms:
To see if the model has semantic understanding of the colour and term encoding that we use, we ask special queries to change the colour of doors/windows and rooms like kitchen, etc.
Arrange Rooms:
To see if the model has spatial understanding, we request specific shapes of the room arrangement like L-, C-, or O-shapes, rectangular shapes, and multiple separated buildings.
For each test prompt and diffusion model in $\mathcal{B}$, $\mathcal{R}$, $\mathcal{SR}$, $\mathcal{SE}$, $\mathcal{SRE}$ we generate 10 sample images. In order to further improve results of all generations, we also added the following negative prompt to each iteration: ”gradients, blurry, fuzzy borders, reflections, lighting”. For each sample image we evaluated manually whether or not (1 or 0) the feature requested in the prompt was addressed in the generated image. This process resulted in 2.150 samples images that we evaluated.
3.2 Discussion
Table 1 shows the results of our evaluation. The baseline out-of-the-box diffusion model $\mathcal{B}$ without fine tuning performs worst. Only 6 % of all generated floor plans were fully valid, while most floor plans were faulty in some aspect like that rooms are missing walls, symbols are partially un-recognizable or room types are not clear (see Fig. 8). Ignoring these flaws, the model generates out-of-the-box something of a floor plan with some symbols for windows and doors and rooms recognizable. This indicates that Stable Diffusion knows the style of floor plans, but is not able to recreate fully.
Table 1: Quantitative results of the experimental validation of the four trained models: ##\## Number of Experiments; $*{}^{*}$ Ignoring full validity
[…]
Valid Plan 410 6 % 63 % 57 % 90 % 66 %
Not Overfitted 20 80 % 95 % 80 % 95 % 100 %
Quantify Obj. 40 65 % 68 % 50 % 55 % 55 %
Quantify Rooms 80 19 % 15 % 58 % 65 % 45 %
Count of Obj. 60 3 % 5 % 2 % 2 % 0 %
Count of Rooms 120 10 % 10 % 12 % 7 % 10 %
Remove Obj. 20 15 % 0 % 20 % 40 % 30 %
Remove Room 30 0 % 0 % 10 % 0 % 13 %
Recoloured Obj. 30 3 % 0 % 0 % 0 % 0 %
Recolour Room 30 0 % 0 % 3 % 0 % 0 %
Arrange Rooms 50 24 % 62 % 34 % 56 % 30 %
Based on this out-of-the-box experience, one may get the impression that Stable Diffusion is unable to generate useful floor plans. Using fine-tuning changes that drastically as visible in the valid plan category. The tuned models mostly manage to display a full building instead of partial cut-outs. There is also recurring symbology that avoids visual artifacts and errors that render a plan difficult to interpret or even entirely unreadable. The model with the best performance here is $\mathcal{SE}$. It is the most visually simple model, without room colouring or complex door and window symbols. Demonstrating that the Simplify strategy helps generating valid models.
The overfitting test results show that the fine-tuned models are not overfitted. Even the base-line models do not always process the overfitting query correctly and all tuned models are in the same margin of error.
Model $\mathcal{SE}$ with semantic element encoding was also the model that performing best to the request to remove specific elements. Most other models perform bad in this category. This shows that here our semantic Encoding strategy helps the diffusion model to learn what windows and doors are. That this is not strictly associated to the colour is evidenced with the results for the recolouring prompts, which usually failed.
Similarly, the results of the count prompts show that the models simply are not yet capable of counting, even after fine-tuning models with many examples with correct counts. The results for quantitative prompts are significantly better. This may partially be because we only have two categories (few/many) resulting in a 50 % chance that the models get it right. This happens for example for the baseline model $\mathcal{B}$, which usually creates incomplete cut-off plans and, in consequential error, gets all few windows/doors/rooms queries correct. Yet, again the semantic encoding model $\mathcal{SE}$ performs above the random draw chance.
The prompts testing room arrangements like L-, C- or O-shapes work well enough across all models at least in comparison to many other queries. This shows that even the baseline model has some basic understanding of arrangements in the higher diffusion model layers. This still is improved significantly with fine-tuning. Here the models $\mathcal{R}$ and $\mathcal{SE}$ show the highest performance increase. This is likely because the walls and rooms are very uniform in colour, increasing the contrast between inside and outside, which is harder to differentiate for the multicoloured rooms. Which supports our Contrasting strategy.
4 Results and Future Directions
4.1 Results
The results from the experiments show that fine-tuning diffusion models significantly improves validity of the generated floor plans from 6 % of the un-tuned model to 57 %–90 % of the fine-tuned models. Tuning the models directly with semantic element encoding $\mathcal{SE}$ is resulting in the best performance across all test prompts. This is proving the correctness of our three strategies Simplify, Encode and Contrast as this model has the best trade-off between them. It is simple in containing only lines. By semantically encoding doors and windows with colours we keep them in the model without reducing simplicity and creating images with high contrast with only 4 colours. Adding room colours is reducing this contrast, which is why the $\mathcal{SR}$ and $\mathcal{SRE}$ models are performing worse.
Our results also show that diffusion models still have major limitations in generating floor plans to satisfy queried requirements. The best performing queries for quantities do not exceed a 68 % success rate. Queries around counts and recolouring usually fail completely. Examples of common errors are shown in Fig. 8. They can be summarized as:
  • Results show only partial floor plans
  • Requested elements are not contained in the generated result
  • Symbols and colour palettes are generated in incorrect context
  • Contextual layout instructions are ignored (e. g. kitchen left of living room)
  • Requested counts of elements and rooms are only met by chance
These issues lead us to some future research directions that we identified.
Figure 8: Examples of generation errors: (a) overlapping door-like symbols; (b) incorrect symbols and colours; (c) building front-view instead of floor-plan
4.2 Future Directions
4.2.1 Future Directions for image-based Diffusion Models
Our experiments revealed several limitations of diffusion models in generating floor plans leading to several future research directions. The success rates of prompts in our experiments are far from optimal. This is one of the reasons why diffusion models usually generate multiple samples at once and let the user to choose. They do than within seconds, significantly outperforming evolutionary generative design approaches. The analysis of Midjourney queries in ploennigs2022ai showed that only 13.1 % of all prompts are upscaled and then refined over 6 iterations—not minimizing the popularity of the platform. Here good workflows that allow to quickly select and refine queries for computational design are important to enable users to quickly filter out bad samples as shown in our idealized workflow in Fig. 2.
Furthermore, the huge body of ongoing research in diffusion models will lead to improved models soon. An important topic in this fundamental area is their capability to correctly count elements in the generated images Yang_2023_CVPR . But, this still requires that the diffusion model has a good understanding of the semantics of the elements in a floor plan. If it does not know what symbol represents a door it will not be able to count it correctly.
This lack of explicit semantic understanding is main limitation of diffusion models. Our experiments showed that it is possible to tune diffusion models to improve the results, but it is not currently feasible to remove all limitations and requires further research in semantic diffusion model encodings. Related to that is the question of how we can decode the semantic encodings of the generated images. Approaches within the area of diffusion models are restoration models, that for example allow to repair faces in generated images wang2021gfpgan .
In relation to that we have the open research question of transforming generated binary images to Building Information Models that we initially wanted as results of the workflow in Fig. 2. The benefit of our semantic encoding is that our strategies of Simplify, Encode, and Contrast also create cleaner images for BIM restoration approaches like pizarro2022automatic .
4.2.2 Future Directions for BIM-based Diffusion Models
All these future research topics above address the limitation of image-based diffusion models in computational design. What we actually want are BIM models. This leaves us with the final essential question: Why not develop a diffusion model for BIM models and simplify the whole process?
From the introduction section 1.2.4 we can derive some requirements for such a BIM-based diffusion model:
  • The model needs to be multi-scale. Traditionally diffusion models work on different resolutions to repair an image successively. These layers then store different information of the original image.
  • The model needs to be diffusable. We need to be able to ingest noise into the model on each level to be able to train a repair network.
  • The model needs to be semantically labelled. We need to be able to label objects and properties as additional input to the diffusion model.
BIM models easily fulfil the last requirement. Given their nature, they are well structured and semantically rich models. Unfortunately, the first two requirements are not easily applicable to BIM models. We can of course raster a BIM model in different resolutions to generate binary images. But, our results show that we lose the semantic information and the well-structured nature in the process and that it is hard to recreate that.
Our proposal is to research in BIM-based Diffusion Models that keep that semantic information and utilizes the hierarchical nature of BIM models. In a BIM model we have floors, zones, and objects that form a hierarchy. Now the only requirement we have to solve is the diffusability. For that we need a Neural Network representation of the BIM model that we can easily alter. Here we can utilize the fact that at the core of BIM models like IFC lies a graph representation like STEP. Representing BIM models in Graphs was proposed before for generative approaches like gan2022bim ; li2023automated . But, we propose to research diffusion models instead of generative models for generation, as they can be trained with additional tags to generate specific designs.
The closest current neural network representation for this are Graph Neural Networks (GNN) recently developed for e. g. time series forecasting ba2022automated . We propose the research into multi-layer GNN model that represents on each layer the respective information of a BIM model, like which floors exists and which are close to each other, which zones exists within these floors and how they connect via walls. Diffusion approaches for graphs capable of weakening or removing those connections in the graph were recently developed chamberlain2021grand .
The benefit of the approach is that we can directly generate 3D models utilizing the BIM structure. A challenge here are the data requirements for training. As discussed in the beginning, training a diffusion model usually requires huge amounts of data. Here, it may be one option to use algorithmic design approaches to generate training data like we did for the fine-tuning. Alternatively, we as research community may need to start an open-source initiative to collect openBIM datasets at large scale to enable this next level of BIM-based Diffusion Models.
(a) Training
(b) Generation
Figure 9: Workflow for a Semantic BIM-based Diffusion Model
5 Conclusion
In this paper we evaluate the use of diffusion models for computational design at the example of floor plans. We analyse the current state of the art in the field and explain diffusion model approaches. We identify the main limitations and develop a novel approach to improve fine-tuning of diffusion models trough semantic encoding. We analyse the workflows for using the diffusion models qualitatively and evaluate the performance of different diffusion models encodings quantitatively in a large set of experiments.
Our experiments show that current diffusion models already create floor plans and that results can be improved significantly with fine-tuning. Particularly our proposed semantic encoding improved validity of the generated floor plans to 90 %. However, our experiments also unveiled several shortcomings in the current diffusion models of which many track back to the lack of semantic understanding. Therefore, we identified several future directions for research in this new field of computation design that may in future automate many detailed planning steps. It is unclear yet, if we can get there with image-based diffusion models or if we need to start working on BIM-based Diffusion Models to unlock their semantic potential.
6 Declarations
Data Availability Statement
The code for generating the training samples as as well as all validation queries will be open sourced on acceptance of the paper.
Funding
No funding was received for conducting this study.
Competing interests
The authors have no financial or proprietary interests in any material discussed in this article.
Author contributions
JP: Ideation and manuscript, Writing Sec. 1, 2.1–2.4, 4, 5. MB: Use Case Experiments, Writing Sec. 2.5, 3. Both authors reviewed and edited all sections as well as read and approved the final manuscript.